General method for formula of circle given two points on circumference and centre lying on a line y = mx +c

Grade:12th pass

1 Answers

Abhishek Mehta
7 years ago
First make an equation of the line by the given two points on circumference Then take intersection with y=mx+c and get the co ordinates of the centre of circleThen find radius and use the formula for circle equation (X-A)^2 +(Y-B)^2 = R^2 Where (A,B) are the co ordinates of centre and R is radius